In a box of 40 cookies, 24 of the cookies were round and 20 of them were made of chocolate. If 12 cookies were neither round nor made of chocolate, how many round chocolate cookies were in the box?
(A)4 (B) 8 (C) 16 (D) 20 (E) 28

Answer by richwmiller(17219)

40-12=28
24+20=44
44-28=16
there were 16 round chocolate cookies
there were 8 other round and 4 other chocolate
and 12 neither round nor chocolate
16+8+4+12=40